What is the result of subtracting the second equation from the first?
-2x+y=0 and -7x+3y=2

Answer by Theo(13342)

your equations are:

-2x + y = 0
-7x + 3y = 2

when you subtract the second equation from the first, you are subtracting like terms from like terms.

-2x - (-7x) = -2x + 7x = 5x
y - 3y = -2y
0 - 2 = -2

the result is 5x -2y = - 2

if a like terms is missing from one of the equation, then substitute the missing term for 0 * the variable that's missing.

example:

y = 0
-7x + 3y = 2

subtracting the second equation from the first, you would do the following.
replace the missing term with 0 * the variable used in that term.

you would get:

0x + y = 0
-7x + 3y = 2

subtract the second equation from the first to get:

7x - 2y = -2

0x - (-7x) is equal to 0x + 7x = 7x
y - 3y = -2y
0 - 2 = -2